find The number of outcomes in the sample space for the gender of the children if a family has 5 children
Answer #1

There are two possible outcomes for the gender i.e. male and female

There are 2 possibility for every child.

Hence total number of outcomes = 2*2*2*2*2 = 2^5 = 32

Ans: 32